# A Lightweight Transformer for Pain Recognition from Brain Activity

### 13 Apr 2026 v1

@@ Abstract
Pain is a multifaceted and widespread phenomenon with substantial clinical and societal burden, making reliable automated assessment a critical objective. This paper presents a lightweight transformer architecture that fuses multiple fNIRS representations through a unified tokenization mechanism, enabling joint modeling of complementary signal views without requiring modality-specific adaptations or increasing architectural complexity. The proposed token-mixing strategy preserves spatial, temporal, and time-frequency characteristics by projecting heterogeneous inputs onto a shared latent representation, using a structured segmentation scheme to control the granularity of local aggregation and global interaction. The model is evaluated on the AI4Pain dataset using stacked raw waveform and power spectral density representations of fNIRS inputs. Experimental results demonstrate competitive pain recognition performance while remaining computationally compact, making the approach suitable for real-time inference on both GPU and CPU hardware.
@@ I Introduction
Pain serves as an evolutionary function by signaling potential tissue damage or disease to the organism, thereby helping preserve physiological integrity [santiago_2022]. It is inherently subjective and multidimensional, involving nociceptive, sensory, affective, and cognitive components [marchand_2024]. Due to its high prevalence and substantial societal burden, pain has been described as a “silent public health crisis ”[katzman_gallagher_2024], emphasizing the need for systematic and reliable assessment. Opioid analgesics remain the most frequently prescribed treatment for pain [kaye_jones_2017], despite their well-documented risks of addiction, overdose [stampas_pedroza_2020], and side effects such as sedation, depression, anxiety, and nausea that severely impair daily functioning and quality of life [benyamin_trescot_2008]. The subjective and complex nature of pain continues to challenge accurate assessment in both clinical practice and research. Reliance on self-reported measures limits treatment precision and contributes to overprescription and misuse of analgesics.
@@
These difficulties are particularly pronounced in medically unstable patients or those unable to communicate effectively [puntilo_staannard_2022]. Pain in critically ill adults is often under-recognized and inadequately managed, mainly due to the scarcity of objective, standardized assessment tools [meehan_mcrae_1995]. Cancer-related pain is especially common, particularly in advanced stages of the disease [snijders_brom_2022]. In addition, pain directly affects attention, functional performance, and cognitive processing speed [khera_rangasamy_2021]. For instance, Li et al. [li_lyu_2025] showed that pain disrupts cognitive control by impairing expectancy processing and decision formation. Moore et al. [moore_meints_2019] demonstrated that both chronic and experimentally induced pain alter attentional performance in a task-dependent manner.
@@
Pain assessment strategies span multiple modalities, each with distinct advantages and limitations. Self-report instruments, such as numerical rating scales and standardized questionnaires, remain the clinical gold standard because they directly capture the subjective experience of pain. However, their reliability is limited in patients who cannot communicate effectively, including critically ill, sedated, cognitively impaired, or pediatric populations. In such cases, behavioral observation methods based on facial expressions, vocalizations, and body movements are used to infer pain levels [rojas_brown_2023]. Physiological signals, including electrocardiography and skin conductance, have also been explored as objective indicators of nociceptive responses [gkikas_tsiknakis_slr_2023]. However, these measures often reflect general stress or arousal rather than pain-specific neural activity, motivating the use of neuroimaging approaches capable of directly capturing cortical processes associated with pain perception.
@@
Functional near-infrared spectroscopy (fNIRS) has emerged as a promising noninvasive neuroimaging modality for capturing cortical activity related to pain. It monitors changes in cerebral hemodynamics and oxygenation by simultaneously measuring concentrations of oxygenated hemoglobin (HbO) and deoxygenated hemoglobin (HbR) in the cortex [rojas_huang_2016]. Studies have consistently shown that noxious stimuli elicit distinct oxygenation changes across multiple cortical regions in both healthy subjects and clinical populations [rojas_liao_2019].
@@
This work explores brain activity using fNIRS, evaluating two representations: the original 1D multichannel waveforms and their time-frequency spectrograms. We introduce a lightweight transformer-based architecture that performs multi-stream, multi-representation fusion through a specialized token-mixing mechanism, achieving strong performance while maintaining computational efficiency.
@@ II Related Work
Over the last decade, research on automatic pain assessment has grown rapidly. This progress has been accompanied by a transition from classical image- and signal-processing pipelines to deep learning-based approaches, although conventional methods still perform well, especially when data are limited [gkikas_phd_thesis_2025]. Most existing methods focus on video, aiming to capture behavioral manifestations of pain through facial expressions, body motion, and other observable cues, and they employ a broad range of learning strategies [huang_dong_2022, bargshady_hirachan_2024, gkikas_tsiknakis_embc, gkikas_tsiknakis_thermal_2024]. Although video-based approaches dominate the field, an increasing number of studies have also investigated biosignal-driven pain recognition, using physiological markers from electrocardiography [gkikas_chatzaki_2022, gkikas_chatzaki_2023], electromyography [thiam_bellmann_kestler_2019, patil_patil_2024, pavlidou_tsiknakis_2025], electrodermal activity (EDA) [ji_zhao_li_2023, phan_iyortsuun_2023, lu_ozek_kamarthi_2023, li_luo_2024, aziz_joseph_2025, gkikas_kyprakis_eda_2025], and brain activity measured with functional near-infrared spectroscopy [rojas_huang_2016, rojas_liao_2019, rojas_romero_2021, khan_sousani_2024, rojas_joseph_bargshady_2024, bargshady_aziz_2025]. Several pain studies have employed fNIRS in conjunction with machine learning to extract informative features and discriminate between pain conditions. In [rojas_huang_2017_b], a bag-of-words representation combined with a K-NN classifier on time–frequency features outperformed using time- or frequency-domain features alone. In contrast, [rojas_juang_2019] reported the best results by combining time- and frequency-domain features with a Gaussian SVM, while [rojas_romero_2021] used raw fNIRS signals with a two-layer BiLSTM, achieving $90.60\%$ accuracy in a multi-class classification task. Finally, [rojas_joseph_2024] proposed a hybrid CNN–LSTM architecture to capture spatio-temporal patterns in fNIRS and reported high performance. Multimodal pain recognition has also gained momentum, with evidence that combining behavioral and physiological information can improve robustness and predictive performance [farmani_bargshady_2025, gkikas_tachos_2024, khan_chetty_2026]. In [shi_chikhaoui_2022], the authors fused features extracted from ECG and EDA and applied tree-based classifiers for pain recognition. Similarly, [jerritta_murugappan_2022] combined wavelet-based features from ECG and EMG and used a k-nearest neighbor classifier. Beyond biosignals, [liang_luo_2025] integrated facial and voice features in pediatric postoperative settings, reporting strong performance and further underscoring the value of multimodal modeling.
@@ III Methodology
This section details the preprocessing pipeline for fNIRS signals, describes the augmentation and regularization strategies used for pain recognition, and presents the proposed transformer model, along with its underlying tokenization approach.
@@ III-A Pre-processing
In the preprocessing stage, the fNIRS signals are utilized in their original waveform format (as multichannel 1D vectors) and also as spectrogram representations derived from the raw signals. These spectrograms are computed via short-time power spectral density (PSD) estimation over successive sliding windows, yielding 2D time-frequency representations that are resized to $224\times 224$ pixels. The proposed methodology can handle both filtered and unfiltered raw signals, as it does not rely on hand-crafted feature engineering. However, in cases of heavy noise contamination, a denoising step may be beneficial. In this study, unfiltered signals are employed.
@@ III-B Architecture
The architecture uses the unified token representation produced by the proposed tokenization method, allowing a single model to process inputs with different dimensionalities, structures, or modalities. A set of learnable latent vectors interacts with the input token sequence through cross-attention and is iteratively refined using self-attention and feedforward layers [gkikas_kyprakis_resp_2025]. This design preserves a fixed latent dimensionality while adapting to variations in input resolution, sequence length, and modality. The approach imposes a structured latent bottleneck by partitioning the token sequence into segments and associating a latent vector with each segment. Specifically, for an input tensor:
@@
$D$ denotes the number of input axes, $A_{1},\ldots,A_{D}$ the axis sizes, and $C$ the number of channels per input location. The number of axes $D$ is kept small (typically $D\in\{1,2\}$) by folding modality-specific factors into the channel dimension. For 2D inputs, $D=2$ and $A_{1}\times A_{2}=H\times W$, where $H$ and $W$ denote the height and width, yielding $\mathbf{X}\in\mathbb{R}^{H\times W\times C}$. Multiple image-like representations per sample (e.g., fNIRS spectrograms) are concatenated along the channel dimension, with $C$ representing the total number of channels. For 1D inputs such as fNIRS waveforms, $D=1$ and $\mathbf{X}\in\mathbb{R}^{L\times C}$, where $L$ is the temporal length and $C$ the number of signal channels.
@@
where $\mathbf{s}=[s_{1},\ldots,s_{K}]^{\top}$ with $\{s_{k}\}_{k=1}^{K}$ spanning $1,f_{\max}/2]$. After flattening, data channels and positional features are concatenated to form the token matrix:
@@
and $\mathbf{Q}_{h}=\mathbf{Q}\mathbf{W}^{Q}_{h}$, with analogous projections for keys and values. Cross-attention and self-attention employ the same mechanism, differing only in the context:
@@
The architecture maintains a structured set of latent states by partitioning the token sequence into $S$ contiguous segments, assigning one latent state to each segment. Tokens are formed as described, yielding $\mathbf{T}\in\mathbb{R}^{N\times C^{\prime}}$. The sequence is partitioned into $S$ segments of length $n_{s}=\lceil N/S\rceil$, padded to $\tilde{N}=Sn_{s}$, and reshaped as:
@@
The tokens of segment $s$ are denoted $\tilde{\mathbf{T}}_{s}\in\mathbb{R}^{n_{s}\times C^{\prime}}$. Segment states are initialized at runtime by replicating a shared learnable vector $\boldsymbol{\ell}_{\mathrm{init}}\in\mathbb{R}^{d}$ across segments. These states are not independent parameters; segment-specific representations emerge from local attention updates. Each layer updates the latent states via attention-based aggregation followed by mixing, with pre-normalization, residual connections, and a position-wise feedforward network. Specifically, each segment state aggregates information from its assigned tokens through local cross-attention:
@@
where $\mathbf{e}^{(\ell)}_{s}\in\mathbb{R}^{d}$ and $\mathrm{Attn}(\mathbf{L},\mathbf{C})$ denotes a pre-normalized multi-head attention block with queries from $\mathbf{L}$ and key-values from $\mathbf{C}$, followed by a feedforward transformation. Stacking segment states yields $\mathbf{E}^{(\ell)}\in\mathbb{R}^{S\times d}$. Inter-segment communication occurs via self-attention on the segment states:
@@
applied multiple times per layer to enhance latent interactions. Local cross-attention is parallelized by treating segments as batch elements, without changing the computation. After $L$ layers, the final segment states $\mathbf{E}^{(L)}$ provides the representation for prediction. Figure 1 illustrates the proposed architecture in a multimodal setup, fusing fNIRS waveforms and their spectrograms. Model hyperparameters for all experiments are summarized in Table I, while Table II presents the training setup.
Figure 1: Overview of the proposed tokenization framework, using videos and fNIRS.
@@ IV Experimental Evaluation & Results
This study utilizes the AI4Pain dataset [ai4pain_2024, rojas_hirachan_2023], comprising synchronized facial videos and fNIRS recordings from $65$ participants. The dataset was gathered at the Human–Machine Interface Laboratory at the University of Canberra, Australia, and divided into $41$ subjects for training, $12$ for validation, and $12$ for testing. To induce controlled pain stimuli, transcutaneous electrical nerve stimulation electrodes were placed on the inner forearm and the dorsum of the right hand. Pain threshold denotes the lowest stimulus intensity perceived as painful (low pain), whereas pain tolerance denotes the highest intensity a participant can tolerate before it becomes unbearable (high pain). For the fNIRS modality, $24$ oxygenated hemoglobin (HbO) channels were employed. All experiments use a three-class classification scheme (No Pain, Low Pain, High Pain) and are evaluated under a hold-out strategy on the validation set, reporting macro-averaged accuracy, precision, and F1 score. Comparisons with previous methods on the testing set are detailed in Section V.
@@ IV-A Impact of Latent Segmentation on fNIRS Representations
Table III reports performance and computational characteristics of the proposed framework on fNIRS inputs across different latent segmentation settings, considering both raw waveform and spectrogram-based PSD representations using all $24$ channels. Varying the number of segments reorganizes how tokens aggregate into latent states without changing token dimensionality, primarily influencing the cost of segment-level interactions and the granularity of localized representation learning. For the waveform representation, accuracy remains stable across segment number, increasing from $43.94\%$ in the unsegmented configuration to $47.18\%$ at $32$ segments, with most settings clustering around the mid-$46\%$ range. Precision and F1 exhibit similar consistency, indicating that segmentation does not introduce unstable optimization behavior. Computationally, waveform inputs remain lightweight, with GFLOPs increasing gradually as the number of segments grows, while GPU latency remains nearly constant (approximately 13131414 ms) across all settings, showing that finer latent structuring improves representational capacity without sacrificing inference speed. For the PSD representation, performance is consistently higher than the waveform across all segment counts. Accuracy remains above $49\%$ throughout, peaking at $52.31\%$ with $2$ segments and remaining strong even as segmentation increases. Precision and F1 follow similar trends, reinforcing that the time–frequency structure provides more discriminative information about pain-related cortical dynamics. Although PSD incurs higher GFLOPs due to higher-dimensional tokenization, GPU latency remains comparable to waveform, indicating efficient scaling with richer representations. Figure 2 presents mean accuracy averaged across waveform and spectrogram representations, alongside GFLOPs and GPU latency for the latent segmentation settings reported in Table III. As the number of segments increases, the computational cost rises gradually, while recognition accuracy remains largely stable, with limited sensitivity to the segmentation level. Segment numbers between $2$ and $16$ yield lower GFLOPs and reduced GPU latency than the unsegmented global representation, while maintaining comparable performance.

### 23 Apr 2026 v2

@@ Abstract
Accurate and reliable pain assessment is an important clinical challenge, especially for patients who cannot communicate their pain directly. In this study, we present a lightweight transformer architecture for pain recognition that leverages multiple fNIRS signal representations and a unified tokenization framework. The proposed architecture enables the integration of heterogeneous inputs in a shared latent space, without requiring separate processing branches or additional architectural complexity. It further employs a structured latent segmentation strategy to capture both local aggregation and global interaction across spatial, temporal, and time–frequency features. Evaluated on the AI4Pain dataset, combining raw waveform and power spectral density fNIRS inputs, the proposed model achieves strong recognition performance across a wide range of computational configurations and, at the same time, maintains efficient inference on both GPU and CPU hardware.
@@ I Introduction
Pain plays a fundamental protective role by alerting the body to possible tissue damage or disease, and supporting physiological integrity [santiago_2022]. By nature, pain is a subjective and multidimensional experience that includes nociceptive, sensory, affective, and cognitive components [marchand_2024]. Pain has been characterized as a “silent public health crisis ”[katzman_gallagher_2024] because of the high prevalence and its impact on society. Opioid analgesics remain the most frequently prescribed treatment for pain [kaye_jones_2017], despite their well-documented risks of addiction, overdose [stampas_pedroza_2020], and side effects such as sedation, depression, anxiety, and nausea that severely impair daily functioning and quality of life [benyamin_trescot_2008]. The subjective and multifaceted nature of pain complicates the accurate assessment in both clinical practice and research. The dependence on self-reported measurement can reduce treatment accuracy and also may contribute to the overprescription and misuse of analgesics [meehan_mcrae_1995].
@@
These challenges are especially pronounced in patients who cannot communicate effectively or directly [puntilo_staannard_2022]. In critically ill adults, pain is often underrecognized and insufficiently treated, because the objective and standardized assessment tools are limited for this category of patients[meehan_mcrae_1995]. Cancer-related pain is especially common, particularly in advanced stages of the disease [snijders_brom_2022]. In addition, pain directly affects attention, functional performance, and cognitive processing speed [khera_rangasamy_2021]. For instance, Li et al. [li_lyu_2025] showed that pain disrupts cognitive control by impairing expectancy processing and decision formation. Moore et al. [moore_meints_2019] demonstrated that both chronic and experimentally induced pain alter attentional performance in a task-dependent manner.
@@
Pain assessment strategies span multiple modalities, each with distinct advantages and limitations. Self-report methods, such as numerical rating scales and standardized questionnaires, remain the clinical gold standard. They reflect the patient’s own experience of pain, however, they when patients cannot communicate effectively, eg., critically ill, sedated, cognitively impaired, or pediatric populations they become less reliable. In these settings, clinicians usually rely on behavioral observation methods, such as facial expressions, vocalizations, and body movements, to estimate pain levels [rojas_brown_2023]. In addition, physiological signals, including electrocardiography and skin conductance, have also been explored as more accurate indicators of nociceptive responses [gkikas_tsiknakis_slr_2023]. However, these measures often capture broader stress or arousal responses rather than neural activity specific to pain. At the same time, developing effective analytical frameworks for the continuous monitoring and assessment of such data remains a wider challenge across domains [chatziadam_dimitraidis_2020].
@@
Functional near-infrared spectroscopy (fNIRS) has emerged as a promising noninvasive neuroimaging modality for capturing cortical activity related to pain. It monitors changes in cerebral hemodynamics and oxygenation by simultaneously measuring concentrations of oxygenated hemoglobin (HbO) and deoxygenated hemoglobin (HbR) in the cortex [rojas_huang_2016]. Broadly, brain activity signals have been shown to reflect a range of affective and cognitive states [gkikas_cruz_eeite_cwl_2026]. Studies have consistently shown that noxious stimuli elicit distinct oxygenation changes across multiple cortical regions in both healthy subjects and clinical populations [rojas_liao_2019].
@@
This work explores brain activity using fNIRS and evaluates two representations: the original 1D multichannel waveforms and their time-frequency spectrograms. We introduce a lightweight transformer-based architecture that performs multi-stream, multi-representation fusion through a specialized token-mixing mechanism, achieving stable performance while maintaining computational efficiency.
@@ II Related Work
Over the last decade, research on automatic pain assessment has grown rapidly. A transition from classical image processing pipelines to deep learning-based approaches has been made, although traditional methods still perform well, especially when data are limited [gkikas_phd_thesis_2025]. Most existing methods focus on video, aiming to capture behavioral manifestations of pain through facial expressions, body motion, and other observable cues, and they employ a broad range of learning strategies [huang_dong_2022, bargshady_hirachan_2024, gkikas_tsiknakis_embc, gkikas_tsiknakis_thermal_2024]. Although video-based approaches dominate the field, an increasing number of studies have also investigated biosignal-driven pain recognition, using physiological markers from electrocardiography [gkikas_chatzaki_2022, gkikas_chatzaki_2023], electromyography [thiam_bellmann_kestler_2019, patil_patil_2024, pavlidou_tsiknakis_2025], electrodermal activity (EDA) [ji_zhao_li_2023, phan_iyortsuun_2023, lu_ozek_kamarthi_2023, li_luo_2024, aziz_joseph_2025, gkikas_kyprakis_eda_2025], and brain activity measured with functional near-infrared spectroscopy [rojas_huang_2016, rojas_liao_2019, rojas_romero_2021, khan_sousani_2024, rojas_joseph_bargshady_2024, bargshady_aziz_2025]. Several pain studies have employed fNIRS in conjunction with machine learning to extract informative features and discriminate between pain conditions. In [rojas_huang_2017_b], a bag-of-words representation combined with a K-NN classifier on time–frequency features outperformed using time- or frequency-domain features alone. In contrast, [rojas_juang_2019] reported the best results by combining time- and frequency-domain features with a Gaussian SVM, while [rojas_romero_2021] used raw fNIRS signals with a two-layer BiLSTM, achieving $90.60\%$ accuracy in a multi-class classification task. Finally, in [rojas_joseph_2024], the authors proposed a hybrid CNN–LSTM architecture to capture spatio-temporal patterns in fNIRS and reported high performance. Multimodal approaches to pain recognition have attracted increasing attention in the research community, since the combination of behavioral and physiological information can improve both robustness and predictive performance [farmani_bargshady_2025, gkikas_tachos_2024, khan_chetty_2026]. In [shi_chikhaoui_2022], the authors fused features extracted from ECG and EDA and applied tree-based classifiers for pain recognition. Similarly, [jerritta_murugappan_2022] combined wavelet-based features from ECG and EMG and used a k-nearest neighbor classifier. Beyond the use of biosignals [liang_luo_2025], a combination of facial and voice features in pediatric postoperative settings has shown strong performance.
Although the literature on automatic pain recognition continues to grow, most existing methods still process different modalities or signal representations through separate feature extraction or modeling pipelines. As a result, only a limited number of studies have explored architectural designs that handle heterogeneous inputs within a single unified framework. The present work addresses this by processing raw waveforms and time-frequency spectrograms through a single tokenization backbone, without representation-specific branches or specialized preprocessing.
@@ III Methodology
This section details the preprocessing pipeline for fNIRS signals and presents the proposed transformer model, along with its underlying tokenization approach.
@@ III-A Pre-processing
In the preprocessing stage, we use the fNIRS signals both in their original waveform form, represented as multichannel 1D sequences, and as spectrograms derived directly from the raw data. Every raw fNIRS channel is converted to a spectrogram by computing its power spectral density (PSD), which measures how the signal energy is distributed across frequencies over short overlapping windows. The result is a 2D time–frequency representation which is then resized to $224\times 224$ pixels. Since the proposed framework does not rely on handcrafted feature extraction, it can be applied to either filtered or unfiltered signals. In this study, we use the unfiltered recordings, allowing the model to learn informative frequency patterns directly from the data while reducing the risk of removing pain-related spectral information through overly aggressive bandpass filtering.
@@ III-B Architecture
The architecture uses the unified token representation produced by the proposed tokenization method, allowing a single model to process inputs with different dimensionalities, structures, or modalities. A set of learnable latent vectors attends to the input tokens through cross-attention and is gradually refined through self-attention and feedforward layers [gkikas_kyprakis_resp_2025]. This method allows the model to keep a fixed latent size while still handling differences in input resolution, sequence length, and modality. The approach imposes a structured latent bottleneck by partitioning the token sequence into segments and associating a latent vector with each segment. Specifically, for an input tensor:
@@
$D$ denotes the number of input axes, $A_{1},\ldots,A_{D}$ the axis sizes, and $C$ the number of channels per input location. The number of axes $D$ is kept small (typically $D\in\{1,2\}$) by folding modality-specific factors into the channel dimension. For 2D inputs, $D=2$ with $A_{1}\times A_{2}=H\times W$, and $\mathbf{X}\in\mathbb{R}^{H\times W\times C}$. If multiple image-like representations exist for a given sample — fNIRS spectrograms being the case here — we stack them along the channel axis, with $C$ counting the total channels. For 1D inputs such as fNIRS waveforms, $D=1$ and $\mathbf{X}\in\mathbb{R}^{L\times C}$, where $L$ is the temporal length and $C$ the number of signal channels.
@@
where $\mathbf{s}=[s_{1},\ldots,s_{K}]^{\top}$ with $\{s_{k}\}_{k=1}^{K}$ spanning $1,f_{\max}/2]$. After flattening, the data channels and positional features are concatenated to create the token matrix:
@@
and $\mathbf{Q}_{h}=\mathbf{Q}\mathbf{W}^{Q}_{h}$, with analogous projections for keys and values. Cross-attention and self-attention employ the same mechanism, but have different contexts:
@@
The architecture maintains a structured set of latent states by partitioning the token sequence into $S$ contiguous segments, assigning one latent state to each segment. The number of latents $M$ reported in Table I corresponds to $S$; it is fixed at $32$ for the fusion experiments but treated as a free parameter in the ablation study. Tokens are formed as described, yielding $\mathbf{T}\in\mathbb{R}^{N\times C^{\prime}}$. The sequence is segmented into $S$ segments of length $n_{s}=\lceil N/S\rceil$, padded to $\tilde{N}=Sn_{s}$, and reshaped as:
@@
The tokens of segment $s$ are denoted $\tilde{\mathbf{T}}_{s}\in\mathbb{R}^{n_{s}\times C^{\prime}}$. At runtime, all segment states start from the same learnable vector $\boldsymbol{\ell}_{\mathrm{init}}\in\mathbb{R}^{d}$ — there are no separate parameters per segment. What differentiates them is purely the result of local attention: each state is updated by attending over its own assigned tokens, and segment-specific structure builds up layer by layer. Each layer applies a cross-attention and is followed by a feedforward layer, with pre-normalization and residual connections. The update for segment $s$ at layer $\ell$ is defined as:
@@
where $\mathbf{e}^{(\ell)}_{s}\in\mathbb{R}^{d}$ and $\mathrm{Attn}(\mathbf{L},\mathbf{C})$ denotes a pre-normalized multi-head attention block with queries from $\mathbf{L}$ and key-values from $\mathbf{C}$, followed by a feedforward layer. By stacking segment states results to $\mathbf{E}^{(\ell)}\in\mathbb{R}^{S\times d}$. Inter-segment communication is happening by applying self-attention on the segment states:
@@
applied multiple times per layer to enhance the interactions between latents. Local cross-attention is parallelized by treating segments as batch elements, without changing the computation. After $L$ layers, the final segment, $\mathbf{E}^{(L)}$, provides the representation for prediction. Figure 1 illustrates the proposed architecture in a multimodal setup, fusing fNIRS waveforms and their spectrograms. Model hyperparameters for all experiments are summarized in Table I, while Table II presents the training setup.
Figure 1: Overview of the proposed tokenization framework, using the two fNIRS representations.
@@ IV Experimental Evaluation & Results
This study uses the AI4Pain dataset [ai4pain_2024, rojas_hirachan_2023], which comprises facial videos and fNIRS recordings from $65$ participants. The data were recorded at the Human–Machine Interface Laboratory at the University of Canberra, Australia, and divided into $41$ subjects for training, $12$ for validation, and $12$ for testing. Pain was induced using transcutaneous electrical nerve stimulation, with electrodes placed on the inner forearm and the back of the right hand. Pain threshold denotes the lowest stimulus intensity perceived as painful (low pain), whereas pain tolerance denotes the highest intensity a participant can tolerate before it becomes unbearable (high pain). For the fNIRS modality, $24$ oxygenated hemoglobin (HbO) channels were employed. All experiments use a three-class classification scheme (No Pain, Low Pain, High Pain) and are evaluated under a hold-out strategy on the validation set, reporting macro-averaged accuracy, precision, and F1 score. Comparisons with previous methods on the testing set are detailed in Section V.
@@ IV-A Impact of Latent Segmentation on fNIRS Representations
Table III reports performance and computational characteristics of the proposed framework on fNIRS inputs across different latent segmentation settings, considering both raw waveform and spectrogram-based PSD representations using all the available $24$ channels. Changing the number of segments affects how tokens are grouped into latent states, without changing their dimensionality. This mainly influences the computational cost and the locality of the learned representations.
For the waveform representation, accuracy is largely stable across segment settings, ranging from $43.94\%$ without segmentation to $47.18\%$ at $32$ segments. The low performance of this representation shows that the raw waveform provides limited discriminative information, regardless of the latent organization. From a computational perspective, the waveforms remain a lightweight modality. Specifically, GFLOPs increase gradually with the number of segments, whereas GPU latency stays nearly unchanged at about 13131414 ms across all settings.
Across all segmentation settings, the PSD representation performs better than the raw waveform, with accuracy above $49\%$ and reaches a peak at $52.31\%$ for $S=2$. This indicates that representing the signal in the time–frequency domain provides more informative pain-related hemodynamic information than using the raw waveform. However, the difference may also partly reflect the higher-dimensional tokenization of PSD inputs. Although PSD incurs higher GFLOPs, GPU latency remains comparable to waveform processing — a practically important property for deployment.
Figure 2 presents the mean accuracy, which is the average across waveform and spectrogram representations, alongside GFLOPs and GPU latency for the latent segmentation settings reported in Table III. We observe that as the number of segments increases, the computational cost rises gradually, while recognition accuracy remains largely stable, with limited sensitivity to the segmentation level. Segment numbers between $2$ and $16$ yield lower GFLOPs and reduced GPU latency than the unsegmented global representation, while maintaining comparable performance.